OpenStudy (anonymous):

What is the solution to the equation 7 −3x ≈ 9?

OpenStudy (anonymous):

x = 0.376 x = 0.295 x = −0.295 x = −0.376

OpenStudy (nikato):

your goal is to isolate x on one side. 1. subtract 7 on both sides. 2. divide both sides by -3.

OpenStudy (anonymous):

so the different equals sign doesnt really matter @nikato

OpenStudy (anonymous):

7-3x=9 -3x=9 -7 -3x=2/-3 x=?

OpenStudy (nikato):

no. thats still an equal sides. it just means approxiamtely. because in this problem, you dont have an exact number for your answer. just an estimate

OpenStudy (anonymous):

when i do that i get -0.66666667

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

@lipowell

OpenStudy (nikato):

i got that too. did you copy your question or answer right?

OpenStudy (anonymous):

yeah @nikato

OpenStudy (nikato):

x=-0.667 should be your solution

OpenStudy (anonymous):

i know thats why i'm confused on why it isn't an option

OpenStudy (nikato):

that should be an option. there might be a typo or something. you should jst ask your teacher

OpenStudy (anonymous):

Thank you for your help!! i was an idiot and typed it wrong. I figured the right answer out. @nikato

OpenStudy (nikato):

oh ok. glad you finally found the answer